Mathematics 243 – Algebraic Structures Definitions Quiz 7 – November 9, 2012

Your Name:

A) (5) Define: d = gcd(a, b) for a, b ∈ Z.

B) (5) Let f : Z_{8} → Z_{8} be defined by f([x]) = [6][x] + [1]. Is f one-to-one? Why or why not?